OpenStudy (anonymous):

A snail travels 3 ft every hour. How far does the snail travel in 2 hours?

OpenStudy (anonymous):

For the next like 30 ish minutes. I'll be asking alot of questions for math, just a heads up

OpenStudy (mrnood):

Now REALLY you must do this yourself. This is a simple 'times table' question. If ALL oyur questions are like this then you should concentrate on art !

OpenStudy (anonymous):

Oh I mean't the snail travels 3 and 1/3 ft every hour and how far does the snail travel in 2 and 1/2 hours

OpenStudy (dls):

1 hour-------3 ft 2 hours--------? u cant even answer that?!

OpenStudy (anonymous):

>.< oops. The copy button doesn't copy fractions guys don't get your panties in a twist!

OpenStudy (anonymous):

sheesh.

OpenStudy (anonymous):

Let's try this again. A snail travels 3 1/3 ft every hour. How far does the snail travel in 2 1/2 hours?

OpenStudy (mrnood):

\[3\frac{ 1 }{ 3 } = \frac{ 4 }{ 3 }\] \[2\frac{ 1 }{ 2 } = \frac{ 3 }{ 2 }\] Multiply the 2 fractions and then rationalise it

OpenStudy (mrnood):

Sorry \[2\frac{ 1 }{ 2 } = \frac{ 5 }{ 2 }\]

OpenStudy (mrnood):

OOPS 3 1/3 = 10/3 bad night!